If the area to the left of x in a normal distribution is 0.163, what is the area to the right of x

33. If the area to the left of x in a normal distribution is 0.163, what is the area to the right of x? (Enter an exact number as an integer, fraction, or decimal)
Answer #1

Area to the left of x in a normal distribution is 0.163

Area to the right of x = 1 - area to the left of x

= 1 - 0.163

= 0.837

34)

Sol

Given

Gender self taught in school private total

Female 11 36 27 74

Male 15 26 15 56

Total. 26 62. 42 130

P( musician is female ORsOR is self taught )

P( F OR ST ) = P( F ) + P( ST ) - P( F AND ST )

F= female

ST = self taught

P( F ) = 74 / 130

P( ST ) = 26/130

P ( F AND ST ) = 11/130

P( F OR ST ) = 74/130 + 26/130 - 11/130

= ( 74 +26 - 11 ) / 130

= 89 / 130

= 0.6846153
Answer #2

ANSWER :


Area to the right of X 

= 1 - area to the left of X

= 1 - 0.163

= 0.837 (ANSWER)

answered by: Tulsiram Garg